OpenStudy (anonymous):

A mass M of 3.60E-1 kg slides inside a hoop of radius R=1.30 m with negligible friction. When M is at the top, it has a speed of 5.17 m/s. Calculate size of the force with which the M pushes on the hoop when M is at an angle of 32.0 degrees.

OpenStudy (anonymous):

Amyvincent are u a memorial student?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

lol basically we solving the last question loooooooooooool,i cant get through with this

OpenStudy (anonymous):

physics 1050

OpenStudy (anonymous):

lon capas are so confusing. I even looked up the methods for each question on google and yet my answers still come up incorrect. and yes, physics 1050.

OpenStudy (anonymous):

i have a link i can give u,but after looking at it i can still really get it right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i'll see if i can figure it out, what topics of the questions don't you understand? i might have some of them done.

OpenStudy (anonymous):

i have all of them done except this last one

OpenStudy (anonymous):

i have idea but am still not arriving at the answer,have 4 trials left

OpenStudy (anonymous):

using this method Etop = mv^2 + mgh = 1/2)(.7)(5.75)^2 + .7(9.8)(2.8) = 30.78 J I Etop = E43degrees 30.78 = .5(.7)(v^2) + .7(1.30(1-cos theta)(9.8) v = 8.32m/s Fnet = mv^2/R mg - Fn = mv^2/R Fn = mv^2/r - mgcos43 Fn = .7(8.32)^2/1.4 - .7(9.8)cos43 Fn= 29.59 N

OpenStudy (anonymous):

using this method,am still getting 17.45N

OpenStudy (anonymous):

i got my answer correct.. i'll tell you the formulas i used in order.. mv^2+mgh = (1/2)mv^2+mgh where h = 2R where the first v^2 is the speed given and the right side of the equation speed is what you're trying to find. once you find that speed, fill it in to FN= (mv^2)/R + mgcos\[\theta\]

OpenStudy (anonymous):

sorry the theta should be on the line above but my computer is weird

OpenStudy (anonymous):

what are your givens?

OpenStudy (anonymous):

m=o.36 r=1.30 v=5.17 theta=32 deg

OpenStudy (anonymous):

Am GETTING 12.84 N

OpenStudy (anonymous):

oh! i have gotten it right the problem was here :Fn = mv^2/r - mgcos43 it should be +

OpenStudy (anonymous):

Sorry its 10.386

OpenStudy (anonymous):

i got the answer as 23.44N,it accepted my answer

OpenStudy (anonymous):

thanks vincent

OpenStudy (anonymous):

perfect, would you mind helping me with another one of my questions i don't have done?

OpenStudy (anonymous):

is it the S1 AND S2 pellet? LOL this :When mass M is at the position shown, it is sliding down the inclined part of a slide at a speed of 2.19 m/s. The mass stops a distance S2 = 1.85 m along the level part of the slide. The distance S1 = 1.17 m and the angle theta = 29.1 degrees. Calculate the coefficient of kinetic friction for the mass on the surface. ?

OpenStudy (anonymous):

yes that one!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what equations did you use?

OpenStudy (anonymous):